Here are the red blood cell counts (in 10^6 cells per microliter) of a healthy person measured on each of 15 days.

5.4 5.3 5.1 5.2 5.5 5.3 5.4 5.1 5.2 5.2 5.4 5.0 5.3 5.3 5.3

Find a 95% confidence interval estimate of μ, the true mean red blood cell count (in 10^6 cells per microliter) for this person during the period of testing. (Round your answers to three decimal places.)

________ 10^6 cells per microliter to _______ 10^6 cells per microliter

You may need to use the appropriate appendix table or technology to answer this question.
Answer #1

sample mean, xbar = 5.267
sample standard deviation, s = 0.1345
sample size, n = 15
degrees of freedom, df = n - 1 = 14

Given CI level is 95%, hence α = 1 - 0.95 = 0.05
α/2 = 0.05/2 = 0.025, tc = t(α/2, df) = 2.145


CI = (xbar - tc * s/sqrt(n) , xbar + tc * s/sqrt(n))
CI = (5.267 - 2.145 * 0.1345/sqrt(15) , 5.267 + 2.145 * 0.1345/sqrt(15))
CI = (5.193 , 5.341)

5.193 10^6 cells per microliter to 5.341 10^6 cells per microliter